Solve for x. Give exact answer. (using logarithm)
13*(1.5)^(x + 3) = 14

Answer by lwsshak3(11628)

Solve for x. Give exact answer. (using logarithm)
13*(1.5)^(x + 3) = 14
log13+(x+3)log1.5=log14
(x+3)=[log14-log13]/log1.5
using calculator
x+3≈0.1828
x≈0.1828-3≈-2.8172